# Rotatable Antenna Enabled Wireless Communication and Sensing: Opportunities and Challenges

### v2

@@ Abstract
Rotatable antenna (RA) is an emerging technology that offers significant potential to enhance wireless communication and sensing performance by flexibly adjusting the boresight of directional antennas. Specifically, RA can flexibly reconfigure its boresight direction via mechanical or electronic means, thereby improving communication channel conditions and/or enhancing sensing resolution and range. In this article, we first provide an overview of RA, including its promising applications, hardware architectures, and radiation pattern characterization. We then illustrate how RA improves communication performance through interference mitigation, spatial multiplexing, and flexible beamforming, as well as sensing capabilities in terms of coverage, resolution, and multi-target/dimensional sensing. Furthermore, we discuss key design challenges in RA systems, including rotational scanning scheduling, channel estimation/sensing, boresight optimization, and RA configuration. Finally, both experimental and simulation results are provided to validate the performance gains achieved by RA for both communication and sensing. Leveraging its unique capabilities in flexible antenna/array rotation to adapt to various communication/sensing requirements and channel conditions, RA is poised to become a key enabler of future intelligent, resilient, and agile wireless networks.
@@ I Introduction
Different from conventional antenna number scaling paradigms, (non-fixed) flexible antenna architectures have emerged as promising alternatives for enhancing wireless system capacity by dynamically reconfiguring antenna characteristics, such as position, shape, and orientation, in response to spatial channel variations. Among them, fluid antenna system (FAS) [2] and movable antenna (MA) [3] have attracted considerable interest owing to their ability to adjust antenna shape and/or position through various mechanisms. These new architectures enable local antenna movement to reposition antennas for creating more favorable channel conditions. However, their performance in practice is often constrained by limited movement speed and considerable hardware response latency. To unlock additional spatial DoFs, six-dimensional MA (6DMA) extends antenna reconfigurability to both three-dimensional (3D) position and rotation adjustments [4]. While this additional flexibility significantly enhances the adaptability to wireless environment, it requires increased design complexity and implementation cost due to the need to satisfy intricate geometric and mechanical constraints for antenna movement.
@@
Motivated by the above, rotatable antenna (RA) [5] has emerged as a viable solution to enhance spatial DoFs while maintaining a compact aperture and low hardware cost. [^1] Historically, RA is not a new concept. Since mechanically controlled RAs originated in the mid-20th century radar systems, RAs have undergone a transformative evolution. Over time, the use of electronic controls and more efficient motors has significantly enhanced the flexibility and performance of RAs. However, earlier research primarily focused on the collective orientation adjustment of entire antenna arrays, with applications largely for radio navigation and localization. In contrast, the RA considered in this article enables the independent boresight direction adjustment of individual antennas, thus offering greater flexibility and higher performance gains for future wireless communication and sensing applications. By independently adjusting the 3D boresight direction of each antenna via mechanical or electronic means, RA can flexibly rotate the radiation pattern towards desired directions. Benefiting from the flexible directional control in 3D space, RA can achieve significant performance gains in terms of precise beam alignment, extended sensing/communication coverage, and adaptive interference management. Although both RA [5] and 6DMA [4] share common rotational flexibility, their architectures and implementations differ significantly. In particular, unlike the translational movement in MA/6DMA, which typically needs additional space (e.g., sliding tracks or movable area), RA only requires local rotational adjustment, which can be achieved more easily through compact mechanical or integrated electronic design. This thus makes RA a streamlined and scalable flexible-antenna technology, offering great compatibility with existing wireless systems. Consequently, RA provides a cost-effective means to exploit spatial DoFs, especially in space and cost constrained scenarios such as IoT nodes and wearable devices.
@@
Moreover, given its unique ability to dynamically adjust the antenna boresight in 3D space, RA technology enables a fundamental paradigm shift in base station (BS) coverage from purely ground-facing service to seamlessly supporting both terrestrial and aerial users/targets. As illustrated in Figure 1, RA opens new opportunities for a wide range of future wireless applications, highlighted as follows.
  • Machine-type communication (MTC): Given the demand for massive device connectivity and ultra-reliable low-latency communication, MTC has become a fundamental pillar of future wireless networks, particularly for emerging applications such as smart cities and industrial automation [6]. However, meeting this demand in practice remains challenging, especially considering the dense node deployment and small-payload devices. In this context, the RA system can adjust antenna’s boresight direction based on the spatial distribution of MTC devices, enabling more effective utilization of the spatial DoFs to enhance signal quality and suppress interference. Furthermore, RA provides a lightweight and scalable solution for MTC devices to deliver high-performance communication and sensing without necessitating large-scale antenna arrays.
  • Space-air-ground integrated network (SAGIN): By seamlessly integrating the space, air, and ground segments, SAGIN aims to deliver global coverage and low-latency communication for future 6G systems [7]. However, conventional down-tilted BSs struggle to support aerial and spaceborne nodes, and their fixed antenna/array orientations further impede the maintenance of reliable links for the dynamic topologies created by mobile aerial and space platforms. To overcome these challenges, RA arrays can be co-deployed at existing BSs to extend communication/sensing coverage across a wider 3D space, thereby enhancing connectivity and tracking capability of the airborne and space segments. Alternatively, RA arrays can be deployed on the aerial or spaceborne nodes to enable flexible beam control, allowing them to adaptively track and maintain high-quality links with ground nodes.
  • Indoor sensing and localization: Indoor sensing and localization are critical for emerging intelligent applications such as smart homes and robotic navigation. Unlike outdoor environments, indoor scenarios usually suffer from severe line-of-sight (LoS) blockage caused by walls, furniture, and other obstacles, which significantly degrades sensing and localization accuracy. This limitation, however, can be alleviated by employing RAs, which rotate their radiation patterns to enhance the directional gain towards LoS reflected paths. Beyond localization, the reconfigurability of RAs is advantageous for simultaneous localization and mapping (SLAM) in robotic networks [8]. By adaptively adjusting the boresight directions, the RA system can efficiently construct channel knowledge maps to facilitate environment-aware communications and sensing with improved accuracy and robustness.
  • Integrated sensing and communication (ISAC): ISAC serves as a key enabler of 6G networks by integrating high-rate communications and high-precision sensing into shared infrastructure and spectrum [9]. However, a key challenge arises when the spatial directions of communication users and sensing targets differ, leading to conflicts in beam steering and resource allocation. To address this issue, RA offers a promising solution by enabling independent 3D boresight control across different antennas in the array. Specifically, RAs can be directed towards different sensing regions to enhance angular resolution and/or aligned with different communication users to improve link quality. Furthermore, by adaptively allocating antenna resources based on the spatial distribution of users and targets, RA can also enhance coverage efficiency in cluttered or dynamic environments.
In light of the above new opportunities, this article aims to provide a comprehensive overview of RA-enabled wireless communication and sensing, with an emphasis on the fundamental principles, hardware architectures, key performance advantages, and practical challenges. Through both experimental and simulation results, we also demonstrate how RA technology can significantly enhance communication and sensing performance of wireless systems under practical constraints.
@@ II Hardware Architecture and Radiation Pattern Characterization
Figure 2 illustrates various hardware architectures for implementing RAs. In addition to the communication module (including radio frequency chain, baseband processing unit, etc.) present in conventional fixed-antenna architectures, RA is incorporated with a dedicated control module for rotating the boresight direction of each antenna or antenna array. Specifically, this boresight adjustment can be achieved through either mechanical or electronic mechanisms, as detailed below.
@@
  • For mechanically driven RAs, as illustrated in Figure 2(a), the antenna’s boresight direction is adjusted by physically rotating the antenna, which directly changes its pointing direction (or equivalently, its orientation) in 3D space. A typical implementation involves mounting the RA on a servo-motor-controlled platform, where the platform’s rotation directly changes the antenna’s orientation in 3D space [10, 11]. Alternatively, mechanically driven RAs can be realized using micro-electromechanical system (MEMS) technology to facilitate miniaturization and simplify deployment [12]. In such systems, torsion hinges are commonly employed to enable high-precision and robust rotational control of individual antennas or an entire array, offering advantages such as compact size, rapid actuation, and low power consumption.
@@
  • For electronically driven RAs, as illustrated in Figure 2(b), the RA’s orientation remains fixed, while its boresight direction is dynamically reconfigured using electronic techniques. For example, for a multi-feed antenna, its main lobe can be rotated in different directions by activating different feed points on a directional radiator [13]. An alternative approach involves the use of parasitic elements with electronic tuning. By tuning the loads (e.g., varactor diodes or PIN diodes) of these parasitic elements, the direction of the radiation pattern can be rotated accordingly [14]. However, these techniques mainly offer discrete and predefined radiation directions. To further enable continuous boresight direction adjustment, more advanced tunable materials (e.g., liquid crystal) can be used.
@@
In general, mechanically driven methods offer a broader control range for boresight adjustment, while electronically driven methods provide much shorter response times and better compatibility with existing wireless platforms. To harness the benefits of both approaches, co-designed RA architectures that integrate both driving mechanisms can be adopted to enable wide-angle and rapid reconfiguration of the radiation pattern, as depicted in Figure 2(c).
@@
In addition to antenna-wise rotation, mechanically driven RA systems can also employ array-level rotation, as illustrated in Figure 2(d). Independently rotating each antenna in an array allows for finer-grained and more flexible boresight reconfiguration. However, this comes at the cost of increased control overhead and design complexity compared to rotating the entire array as a unit. Consequently, dual-scale architectures with both antenna-wise and array-wise rotation can be adopted to further enhance rotation resolution and range, as shown in Figure 2(e). Furthermore, to enhance compatibility with existing BS architectures, hybrid configurations that integrate RAs into fixed-antenna BSs can be adopted, as depicted in Figure 2(f). In this setup, the sector antennas retain their inherent capability to serve terrestrial users, while the co-deployed antenna/array-level RAs extend coverage to serve both terrestrial and aerial users/targets. This design enhances the system’s overall connectivity and coverage performance, particularly benefiting emerging low-altitude economy with rich UAV applications.
@@
Besides hardware architecture, a fundamental aspect of RA is the characterization of how its radiation pattern is manipulated during operation. As shown in Figure 2, both MA and RA preserve the intrinsic shape of their antenna radiation patterns during reconfiguration, but they manipulate the pattern through fundamentally different mechanisms. Specifically, MA translates the radiation pattern by relocating the antenna’s position, whereas RA rotates its radiation pattern by adjusting the antenna’s or array’s boresight direction, thereby enabling angular diversity to enhance communication and sensing performance. Nevertheless, the radiation pattern rotation flexibility in practice depends on implementation feasibility and environmental constraints. For example, high rotation resolution demands either a high-precision actuator for mechanically driven RAs or a large number of feasible switching or excitation states for electronically driven RAs, both of which increase hardware cost. Moreover, the rotation range of RA also depends on the adjustable range of the control system and the surrounding environment. By selecting an appropriate control accuracy and range, the radiation pattern of RA can be effectively rotated to allocate signal power towards desired directions, thereby enhancing directional gain and suppressing interference.
@@ III-A Wireless Communication
  • Spatial multiplexing: In rich scattering environments, MIMO systems enhance channel capacity by spatially multiplexing multiple data streams for simultaneous transmission. The achievable spatial multiplexing gain hinges on rich scattering which enables receive antennas to effectively distinguish signals from different transmit antennas. However, this channel condition may not be achieved in practice when there exists a dominant LoS component or strong reflected paths, especially in near-field scenarios, resulting in a reduced effective channel rank and, consequently, degraded achievable multiplexing gain. To alleviate such performance degradation, RA provides a new means to create subchannel orthogonality and thereby improve the channel conditions [15]. Through dynamic boresight adjustment, RA arrays enhance the angular separation among signal paths, which improves spatial resolvability and facilitates concurrent transmissions of multiple data streams in both point-to-point multi-user communication systems.
@@ III-B Wireless Sensing
IV Design Challenges: Unlocking the Full Potential of RA
@@
IV-A Rotational Scanning Scheduling
@@
Generally, rotational scanning scheduling poses several key challenges. First, the design of scanning scheduling must account for multiple factors, including the scanning parameters (e.g., range, frequency, pattern) as well as the coordination with other signal processing tasks to ensure efficient system operation. This makes the scheduling problem inherently intricate. Second, different tasks, such as communication, sensing, navigation, and localization, impose diverse requirements on scanning resolution, latency, and angular precision, thereby necessitating adaptive and context-aware scheduling strategies. To address these challenges, apart from the straightforward round-robin strategy, several effective solutions can be adopted. In cases where prior information is available, such as channel statistics or environmental information, RA can adaptively tailor its scanning pattern and range to reduce redundant scans. On the other hand, when such prior information is unavailable or unreliable, a coarse-to-fine scanning strategy can be adopted to effectively balance scanning resolution and latency. To further reduce the scanning overhead, multi-RA cooperative scheduling designs can be employed. In multi-RA systems, dividing the antennas into groups for performing different scanning tasks can help reduce the workload per RA while maintaining full coverage, which is particularly useful in large-scale deployments or distributed sensing networks.
IV-B Channel Estimation/Sensing

### v3

@@ Abstract
Rotatable antenna (RA) is an emerging technology that offers significant potential to enhance wireless communication and sensing performance by flexibly adjusting the boresight of directional antennas. Specifically, RA can flexibly reconfigure its boresight direction via mechanical or electronic means, thereby improving communication channel conditions and/or enhancing sensing resolution and range. In this article, we first provide an overview of RA, covering its hardware architectures and radiation pattern characterization. We then illustrate how RA improves communication performance through interference mitigation, spatial multiplexing, and flexible beamforming, as well as sensing capabilities in terms of coverage, resolution, and multi-target/dimensional sensing. Furthermore, we highlight representative applications of RA and discuss key design challenges in RA systems, including rotational scanning scheduling, channel estimation/sensing, boresight optimization, and RA configuration. Finally, both experimental and simulation results are provided to validate the performance gains achieved by RA for both communication and sensing. Leveraging its unique capabilities in flexible antenna/array rotation to adapt to various communication/sensing requirements and channel conditions, RA is poised to become a key enabler of future intelligent, resilient, and agile wireless networks.
@@ I Introduction
Different from conventional antenna number scaling paradigms, (non-fixed) flexible antenna architectures have emerged as promising alternatives for enhancing wireless system capacity by dynamically reconfiguring antenna characteristics, such as position, shape, and orientation, in response to spatial channel variations. Among these architectures, fluid antenna system (FAS) [2] and movable antenna (MA) [3] have attracted considerable interest because of their ability to adjust antenna shape and/or position to create more favorable channel conditions. However, their performance in practice is often constrained by limited movement speed and considerable hardware response latency. To unlock additional spatial DoFs, six-dimensional MA (6DMA) extends antenna reconfigurability to both three-dimensional (3D) position and rotation adjustments [4]. While this additional flexibility significantly enhances the adaptability to the wireless environment, it requires increased design complexity and implementation cost due to the intricate geometric and mechanical constraints for antenna movement.
@@
Motivated by the above, rotatable antenna (RA) [5] has emerged as a viable solution to enhance spatial DoFs while maintaining a compact aperture and low hardware cost. [^1] Historically, RA is not a new concept. Since mechanically controlled RAs originated in the mid-20th century radar systems, RAs have undergone a transformative evolution. Over time, the use of electronic controls and more efficient motors has significantly enhanced the flexibility and performance of RAs. Earlier research primarily focused on the orientation adjustment of entire antenna arrays, with applications largely in radio navigation and localization[6]. In contrast, the RA considered in this article enables independent boresight direction adjustment of individual antennas within an array, thus offering greater flexibility, more intelligence, and higher performance gains for next-generation wireless systems. By independently adjusting the 3D boresight direction of each antenna via mechanical or electronic means, RA can rotate the radiation pattern towards desired directions, thereby achieving significant performance gains in terms of precise beam alignment, extended coverage, and adaptive interference management. Table I summarizes and compares the key characteristics of representative flexible antenna architectures, including FAS, MA, 6DMA, flexible antenna array (FAA) [7], and RA. Although RA [5] can be considered as a special case of 6DMA[4, 8] in terms of movement model by sharing the common rotational flexibility, their practical architectures and implementations differ significantly. In particular, unlike the translational movement in MA/6DMA, which typically needs additional space (e.g., sliding tracks or movable area), RA only requires local rotational adjustment, which can be achieved more easily through compact mechanical or integrated electronic design. This makes RA a streamlined and scalable flexible-antenna solution, offering great compatibility with existing wireless systems as well as enabling efficient deployment in space and cost-constrained scenarios such as IoT nodes and wearable devices.
TABLE I: Comparison of Different Flexible Antenna Architectures
Antenna
Architecture
Hardware Implementation
Reconfigurable Parameter
Deployment
Complexity
System
Overhead
RA [5]
Motors or electronic means
Antenna 3D boresight direction
Very Low Very Low
FAA [7] Flexible materials Array shape/structure Moderate Moderate
FAS [2] Pixel antennas or liquid materials
Antenna shape and 3D position
Moderate Moderate
MA [3] Motors Antenna 3D position Moderate Moderate
6DMA [4] Motors and flexible cables
Antenna 3D position and 3D rotation
Moderate Moderate
@@
Given its unique ability to dynamically adjust the antenna boresight in 3D space, RA technology enables a fundamental paradigm shift in base station (BS) coverage from purely ground-facing service to seamlessly supporting both terrestrial and aerial users/targets. As illustrated in Fig. 1, RA opens new opportunities for a broad range of future wireless applications, which will be further elaborated in Section IV. Motivated by the above, this article aims to provide a comprehensive overview of RA-enabled wireless communication and sensing, with an emphasis on the fundamental principles, hardware architectures, key performance advantages, potential applications, and practical challenges. Through both experimental and simulation results, we also demonstrate how RA technology can significantly enhance communication and sensing performance of wireless systems under practical constraints.
@@ II Hardware Architecture and Radiation Pattern Characterization
Fig. 2 illustrates various hardware architectures for implementing RAs. In addition to the communication module (including radio frequency chain, baseband processing unit, etc.) present in conventional fixed-antenna architectures, RA is incorporated with a dedicated control module for rotating the boresight direction of each antenna or antenna array. Specifically, this boresight adjustment can be achieved through either mechanical or electronic mechanisms, as detailed below.
@@
  • As illustrated in Fig. 2(a), mechanically driven RA employs actuators to physically rotate the antenna orientation, thereby altering its boresight direction. A typical implementation involves mounting the RA on a servo-motor-controlled platform, where the platform’s rotation directly adjusts the antenna’s orientation in 3D space [9, 10]. Alternatively, mechanical adjustment can be achieved using micro-electromechanical system (MEMS) technology to facilitate miniaturization and simplify deployment [11]. In such systems, torsion hinges are commonly employed to enable high-precision and robust rotational control of the antenna, offering advantages including compact size, rapid actuation, and low power consumption.
@@
  • As illustrated in Fig. 2(b), electronically driven RA maintains fixed antenna orientation, while its boresight direction is dynamically reconfigured through electronic techniques. For example, for a multi-feed antenna, its main lobe can be rotated in different directions by activating different feed points on a directional radiator [12]. An alternative approach involves the use of parasitic elements with electronic tuning. By tuning the loads (e.g., varactor diodes, PIN diodes) of these parasitic elements, the direction of the antenna radiation pattern can be rotated accordingly [13]. However, these techniques mainly offer discrete and predefined radiation directions. To further enable continuous boresight direction adjustment, more advanced tunable materials (e.g., liquid crystal [14]) can be used.
@@
In general, mechanically driven methods offer a broader control range for boresight adjustment, while electronically driven methods provide better compatibility with existing wireless platforms. In practice, MEMS-based actuators typically entail a power consumption in the milliwatt range, with response time on the order of microseconds to milliseconds. Meanwhile, electronically driven RAs can typically achieve response time ranging from nanoseconds to milliseconds. To harness the benefits of both approaches, co-designed RA architectures that integrate both control mechanisms can be adopted to enable wide-angle and rapid reconfiguration of the radiation pattern, as depicted in Fig. 2(c).
@@
In addition to antenna-level rotation, mechanically driven RA systems can also incorporate array-level rotation, as illustrated in Fig. 2(d). Independently rotating each antenna in an array allows for finer-grained and more flexible boresight reconfiguration. However, this comes at the cost of increased control overhead and design complexity compared to rotating the entire array as a unit. On the other hand, array-level rotation inherently alters the antenna radiation pattern, which may require additional compensation mechanisms such as antenna rotation calibration to maintain the desired system performance. Consequently, dual-scale architectures with both antenna-level and array-level rotation can be adopted to further enhance rotation resolution and range, as shown in Fig. 2(e). Furthermore, to enhance compatibility with existing BS architectures, hybrid configurations that integrate RAs into fixed-antenna BSs can be adopted, as depicted in Fig.  2(f). In this setup, the sector antennas retain their inherent capability to serve terrestrial users, while the co-deployed antenna/array-level RAs extend coverage to serve both terrestrial and aerial users/targets. This design enhances the system’s overall connectivity and coverage performance, particularly benefiting emerging low-altitude economy with rich UAV applications.
@@
Besides hardware architecture, a fundamental aspect of RA is the characterization of how its radiation pattern is manipulated during operation. As shown in Fig. 2, both MA and RA preserve the intrinsic shape of their antenna radiation patterns during reconfiguration, but they manipulate the pattern through fundamentally different mechanisms. Specifically, MA translates the radiation pattern by relocating the antenna’s position, whereas RA rotates its radiation pattern by adjusting the antenna’s or array’s boresight direction, thereby enabling angular diversity to enhance communication and sensing performance. Nevertheless, the radiation pattern rotation flexibility in practice depends on implementation feasibility and environmental constraints. For example, high rotation resolution demands either a high-precision actuator or a large number of feasible switching or excitation states, both of which incur substantial hardware cost. Moreover, the rotational range of RA also depends on the adjustable range of the control system and the surrounding environment. By selecting an appropriate control accuracy and range, the radiation pattern of RA can be effectively rotated to allocate signal power towards desired directions, thereby enhancing directional gain and suppressing interference.
@@ III-A Wireless Communication
  • Spatial multiplexing: In rich scattering environments, MIMO systems enhance channel capacity by spatially multiplexing multiple data streams for simultaneous transmission. The achievable spatial multiplexing gain hinges on rich scattering which enables receive antennas to effectively distinguish signals from different transmit antennas. However, this channel condition may not be achieved in practice when there exists a dominant line-of-sight (LoS) component or strong reflected paths, especially in near-field scenarios, resulting in a reduced effective channel rank and, consequently, degraded achievable multiplexing gain. To alleviate such performance degradation, RA provides a new means to create subchannel orthogonality and thereby improve the channel conditions. Through dynamic boresight adjustment, RA arrays enhance the angular separation among signal paths, which improves spatial resolvability and facilitates concurrent transmissions of multiple data streams in both point-to-point multi-user communication systems.
@@ III-B Wireless Sensing
IV Potential Applications of RA
Building on the aforementioned performance advantages, RA enables a wide range of promising applications in future wireless systems. In this section, we highlight several representative scenarios where the boresight reconfigurability of RA can be effectively leveraged to meet increasingly stringent requirements in terms of capacity, coverage, spatial resolution, and other key system performance metrics for emerging applications.
  • Machine-type communication (MTC): Given the demand for massive device connectivity and ultra-reliable low-latency communication, MTC has become a fundamental pillar of future wireless networks, particularly for emerging applications such as smart cities and industrial automation. However, meeting this demand in practice remains challenging, especially considering the dense node deployment and small-payload devices. In this context, the RA system can adjust antenna’s boresight direction based on the spatial distribution of MTC devices, enabling more effective utilization of the spatial DoFs to enhance signal quality and suppress interference. Furthermore, RA provides a lightweight and scalable solution for MTC devices to deliver high-performance communication and sensing without necessitating large-scale antenna arrays.
  • Space-air-ground integrated network (SAGIN): By seamlessly integrating the space, air, and ground segments, SAGIN aims to deliver global coverage and low-latency communication for future 6G systems. However, conventional down-tilted BSs struggle to support aerial and spaceborne nodes, and their fixed antenna/array orientations further impede the maintenance of reliable links for the dynamic topologies created by mobile aerial and space platforms. To overcome these challenges, RA arrays can be co-deployed at existing BSs to extend communication/sensing coverage across a wider 3D space, thereby enhancing connectivity and tracking capability of the airborne and space segments. Alternatively, RA arrays can be deployed on the aerial or spaceborne nodes to enable flexible beam control, allowing them to adaptively track and maintain high-quality links with ground nodes.
  • Indoor sensing and localization: Indoor sensing and localization are critical for emerging intelligent applications such as smart homes and robotic navigation. Unlike outdoor environments, indoor scenarios usually suffer from severe LoS blockage caused by walls, furniture, and other obstacles, which significantly degrades sensing and localization accuracy. This limitation, however, can be alleviated by employing RAs, which rotate their radiation patterns to enhance the directional gain towards LoS reflected paths. Beyond localization, the reconfigurability of RAs is advantageous for simultaneous localization and mapping (SLAM) in robotic networks. By adaptively adjusting the boresight directions, the RA system can efficiently construct channel knowledge maps to facilitate environment-aware communications and sensing with improved accuracy and robustness.
  • Integrated sensing and communication (ISAC): ISAC serves as a key enabler of 6G networks by integrating high-rate communications and high-precision sensing into shared infrastructure and spectrum. However, a key challenge arises when the spatial directions of communication users and sensing targets differ, leading to conflicts in beam steering and resource allocation. To address this issue, RA offers a promising solution by enabling independent 3D boresight control across different antennas in the array. Specifically, RAs can be directed towards different sensing regions to enhance angular resolution and/or aligned with different communication users to improve link quality. Furthermore, by adaptively allocating antenna resources based on the spatial distribution of users and targets, RA can also enhance coverage efficiency in cluttered or dynamic environments.
V Design Challenges: Unlocking the Full Potential of RA
@@
V-A Rotational Scanning Scheduling
@@
Generally, rotational scanning scheduling poses several key challenges. First, the design of scanning scheduling must account for multiple factors, including the scanning parameters (e.g., range, frequency, pattern) as well as the coordination with other signal processing tasks to ensure efficient system operation. This makes the scheduling problem inherently intricate. Second, different tasks, such as communication, sensing, navigation, and localization, impose diverse requirements on scanning resolution, latency, and angular precision, thereby necessitating adaptive and context-aware scheduling strategies. To address these challenges, several effective solutions can be adopted. In cases where prior information is available, such as channel statistics or environmental information, RA can adaptively tailor its scanning pattern and range to reduce redundant scans. On the other hand, when such prior information is unavailable or unreliable, a coarse-to-fine scanning strategy can be adopted to effectively balance scanning resolution and latency. To further reduce the scanning overhead, multi-RA cooperative scheduling designs can be employed. In multi-RA systems, dividing the antennas into groups for performing different scanning tasks can help reduce the workload per RA while maintaining full coverage, which is particularly useful in large-scale deployments or distributed sensing networks.
V-B Channel Estimation/Sensing